OpenStudy (sprattfield):

Fan and medal Evaluate the following expression using the values given: Find 2x − 3y − z if x = −2, y = 3, and z = −2 −11 11 −15 15

OpenStudy (1davey29):

Just plug in the values. 2(-2)-3(3)-(-2)=-4-9+2=-11

OpenStudy (otherworldly):

-15